# High-performance Effective Scientific Error-bounded Lossy Compression with Auto-tuned Multi-component Interpolation

### v2

@@ 7.1.3. Experimental configurations and evaluation metrics
  • Compression/decompression speed: Check the compression and decompression speeds of compressors.
@@
  • Rate-PSNR plots: Plot curves for compressors with the bit rate of the compressed data and the decompression PSNR.
@@ 7.2.1. Speeds
To verify our categorization of compressors and examine the compression efficiency of HPEZ, in Table 2 we present the compression and decompression speeds of 6 comparison compressors and HPEZ (under error bound 1e--3, i.e., $10^{-3}$) on the Anvil machine. From the table, we can clearly observe that the high-performance compressors (SZ 3.1, ZFP 0.5.5, and QoZ 1.1) have far better compression speeds than the high-ratio compressors (SPERR, FAZ, and TTHRESH) with the gap of 3$\times$ $\sim$ 10$\times$. Having a speed of around 70% $\sim$ 90% of QoZ 1.1, HPEZ can definitely be regarded as a high-performance compressor, achieving 2$\times$ $\sim$ 6$\times$ performance improvement over SPERR/FAZ, and 4$\times$ $\sim$ 17$\times$ performance improvement over TTHRESH. This relatively high speed ensures the advantages of HPEZ on efficiency-oriented and high-ratio-preferred compression tasks. Figure 10 presents the error bound-compression speed curves of each compressor on the 6 tested datasets ( the x-axis is the negative log10 of the error bounded and the y-axis is the compression speed). Those plots also prove that HPEZ is much more efficient than the high-ratio compressors (SPERR, FAZ, and TTHRESH) and has close performances to SZ3 and QoZ.
@@
Type Dataset SZ ZFP QoZ SPERR FAZ TTH HPEZ 3.1 0.5.5 1.1 0.6 RESH Compression CESM 219 331 215 49 58 10 140 RTM 211 412 191 63 30 18 142 Miranda 163 416 157 35 29 28 140 SCALE 188 191 182 32 61 17 129 JHTDB 140 225 122 33 28 23 105 SegSalt 189 645 201 51 36 13 141 Decompression CESM 661 584 689 92 101 53 513 RTM 786 622 626 124 64 108 510 Miranda 419 946 351 75 60 111 473 SCALE 610 553 567 68 140 53 450 JHTDB 376 425 243 70 59 60 330 SegSalt 592 1060 629 108 65 97 485
@@ 7.2.2. Compression ratios with the same error bounds
Compressing the datasets with the selected compressors under the same error bounds, we list all the compression ratios acquired in Table 3 and 4. Table 3 is a comparison among the high-performance compressors (in which the compression ratio optimization targets are set for QoZ, FAZ, and HPEZ). HPEZ achieves the best compression ratio in all cases. On the SegSalt dataset, HPEZ has a 40% $\sim$ 75% compression ratio improvement over the second-best compressor. On the RTM, Miranda, and JHTDB datasets, HPEZ achieves 20%-45% compression ratio improvements over the second-best. On the CESM-ATM dataset, under the error bound of 1e-3, HPEZ has a compression ratio of about 2.36$\times$ as high as the second-best (SZ3.1). With these considerable improvements, we can assert that HPEZ is the best choice among high-performance compressors regarding optimizing the error-bound-fixed compression ratio.
@@
Dataset $\epsilon$ SZ ZFP QoZ HPEZ Improve 3.1 0.5.5 1.1 (%) RTM 1E-2 1764 62.9 2156 2701 25.3 1E-3 249 26.2 285 395 38.6 1E-4 55.3 14.3 58 71.1 22.6 Miranda 1E-2 574.6 46.6 977 1320 35.1 1E-3 168 25.6 181 258 42.5 1E-4 47.3 14.5 47.7 63.6 33.3 SegSalt 1E-2 856 59.1 1005 1484 47.7 1E-3 140.6 24.9 151 260 72.2 1E-4 38.2 14.9 35.9 61.7 61.5 SCALE 1E-2 167.3 14.5 160 186 11.2 1E-3 40.4 7.8 41.5 52.9 27.5 1E-4 14.1 4.6 13.4 15.4 9.2 JHTDB 1E-2 528.2 22.3 647 838 29.5 1E-3 73.2 9.8 77.8 101 29.8 1E-4 15.8 5 15.9 20.6 29.6 CESM-ATM 1E-2 373 18.2 263 675 81.0 1E-3 64.9 9.6 59.4 153 135.7 1E-4 22.9 5.8 21.7 38.9 69.9
@@
Dataset $\epsilon$ SPERR FAZ TTHRESH HPEZ 0.6 RTM 1E-2 2187 2695 782 2701 1E-3 440 642 71.4 395 1E-4 84.1 119 23.7 71.1 Miranda 1E-2 971.4 996.5 447 1320 1E-3 243.9 263.5 142 258 1E-4 74.5 93.6 55.1 63.6 SegSalt 1E-2 1219.4 1639.6 291 1484 1E-3 228.9 388.9 99.5 260 1E-4 61.3 117.3 28.8 61.7 SCALE 1E-2 103.5 177.9 80.0 186 1E-3 35.5 51.8 18.9 52.9 1E-4 15 16.8 8.4 15.4 JHTDB 1E-2 639.8 726 373 838 1E-3 89.3 90.7 65.1 101 1E-4 19.9 20.2 17.1 20.6 CESM-ATM 1E-2 1221 292 83.5 675 1E-3 150 77.4 20.4 153 1E-4 35 26.3 8.7 38.9
@@ 7.2.3. Compression rate-distortion
Figure 13. Rate-PSNR of HPEZ and high-ratio compressors (HPEZ’s speed is 2x-17x of the high-ratio compressors).
@@ 7.2.4. Parallel data transfer test on the distributed database
Dataset Direction Inter-machine SZ3 ZFP QoZ 1.1 SPERR 0.6 FAZ TTHRESH HPEZ Improve
Speed (GB/s) (%)
CESM-ATM Anvil to Bebop 0.79 $\sim$0.91 1934 3221 1812 1560 1586 7752 1005 35.6
(41TB) Bebop to Anvil 0.95 $\sim$1.19 1614 2695 1553 1522 1544 8560 916 39.8
RTM Anvil to Bebop 0.58 $\sim$1.19 198 362 173 277 494 527 181 -4.8
(14TB) Bebop to Anvil 0.47 $\sim$1.04 189 524 166 296 474 560 182 -9.5
Miranda Anvil to Bebop 0.46 $\sim$1.04 49 84 44 72 87 121 39 11.3
(2TB) Bebop to Anvil 0.54 $\sim$0.82 46 117 49 71 86 120 43 6.5
SCALE-LetKF Anvil to Bebop 0.88 $\sim$0.94 873 1354 820 1037 782 2354 728 7.0
(13TB) Bebop to Anvil 1.05 $\sim$1.15 745 1181 707 1007 670 2002 624 6.8
JHTDB Anvil to Bebop 0.83 $\sim$1.15 567 826 527 645 583 835 417 20.9
(10TB) Bebop to Anvil 0.97 $\sim$1.18 486 707 473 648 574 883 366 22.7
SegSalt Anvil to Bebop 0.63 $\sim$1.18 163 289 174 221 251 393 137 15.9
(8TB) Bebop to Anvil 0.76 $\sim$1.06 167 241 153 213 265 300 132 14.0
@@ 7.2.5. Case study: decompression visualizations
As an example of the effectiveness of the HPEZ compression, in this section, we propose a case study of the compression tasks, featuring and visualizing the decompression outputs from various high-performance compressors. The example data input we choose is the QS field (getting logarithmized in preprocessing) from the SCALE-LetKF dataset, and we compress it with HPEZ and 2 high-performance compressors: QoZ and ZFP (we omit SZ3 in this test because QoZ and SZ3 have close speeds and QoZ has better compression quality than SZ3) under similar compression ratios. The visualizations of 2-D slices from the original data and decompressed data are presented in Figure 15. In this case, among the decompression results with very close compression ratios, the decompression result of HPEZ (Figure 15 (b)) achieves the lowest data distortion with the highest PSNR (56.8). Moreover, regarding the magnified regions in Figure 15, compared to the decompression results of QoZ (Figure 15 (c), PSNR=52.7), HPEZ has better preserved the local data patterns in the original input (Figure 15 (a)). This case is an example to show the strong capability of HPEZ in providing high-quality compression results with high compression speed.
@@
Figure 15. Visualization of SCALE-QS field (logarithmized) and the decompressed data with different compressors.
@@ 7.2.7. Ablation study
Figure 17 shows the rate-PSNR plots on the RTM dataset with QoZ 1.1, HPEZ, and the different accumulations of new design components between them. For example, in Figure 17 (a) representing the results of the RTM dataset, the blue curve shows the rate-distortion of QoZ 1.1, the orange curve shows the rate-distortion of QoZ 1.1 plus the interpolation re-ordering, the green curve is for the orange curve plus the natural cubic spline, and so on, eventually to the complete HPEZ. For the RTM, JHTDB, Miranda, and SegSalt datasets (Figure 17 (a) (c) (d) (f)), analyzing the rate-distortion curves we can easily find that the HPEZ interpolation designs, including interpolation re-ordering (Section 5.4), natural cubic spline (Section 5.2), and multi-dimensional interpolation (Section 5.3) all contribute to the improvement of rate-distortion. Additionally, block-wise interpolation tuning (Section 6.6) also plays an important role in optimizing the compression of their compression. Lastly, the effectiveness of multi-dimensional spline interpolation proved the generalization of Theorem 5.1 on diverse datasets and the integration of multiple interpolation optimization techniques.
In Figure 17 (b) and (e) corresponding to the CESM-ATM and SCALE-LetKF datasets, we can verify the effectiveness of dynamic dimension freezing (Section 6.3) and the Lorenzo predictor (Section 6.5) The orange dashed line in Figure 17 (b) and (e) integrates all the interpolation designs in Section 5. Nevertheless, compared with QoZ 1.1 they have not refined the compression sufficiently. In contrast, the dynamic dimension freezing itself (the red curve in Figure 17 (b) and the green curve in Figure 17 (e)) has solely boosted the rate-distortion to a remarkable extent for those 2 datasets. Furthermore, comparing the red solid line and the green dash-dot line in Figure 17 (e), leveraging the Lorenzo predictor has quite enhanced the compression quality of HPEZ in low-error-bound (i.e. high bit rates) cases.

### v3

@@ 7.1.3. Experimental configurations and evaluation metrics
  • Speeds: Check the compression and decompression speeds of compressors.
@@
  • Rate-PSNR plots: Plot curves for compressors with the bit rate of the compressed data and the decompression data PSNR.
@@ 7.2.1. Speeds
To verify our categorization of compressors and examine the compression efficiency of HPEZ, in Table 2 we present the compression and decompression speeds of 6 comparison compressors and HPEZ (under error bound 1e--3, i.e., $10^{-3}$) on the Anvil machine. From the table, we can clearly observe that the high-performance compressors (SZ 3.1, ZFP 0.5.5, and QoZ 1.1) have far better compression speeds than the high-ratio compressors (SPERR, FAZ, and TTHRESH) with the gap of 3$\times$-10$\times$. Having a speed of around 70% $\sim$ 90% of QoZ 1.1, HPEZ can definitely be regarded as a high-performance compressor, achieving 2$\times$ $\sim$ 6$\times$ performance improvement over SPERR/FAZ, and 4$\times$ $\sim$ 17$\times$ performance improvement over TTHRESH. This relatively high speed ensures the advantages of HPEZ on efficiency-oriented and high-ratio-preferred compression tasks. Figure 10 presents the error bound-compression speed curves of each compressor on the 6 tested datasets ( the x-axis is the negative log10 of the error bounded and the y-axis is the compression speed). Those plots also prove that HPEZ is much more efficient than the high-ratio compressors (SPERR, FAZ, and TTHRESH) and has close performances to SZ3 and QoZ.
@@
Type Dataset SZ 3.1 ZFP 0.5.5 QoZ 1.1 SPERR 0.6 FAZ TTHRESH HPEZ Compression CESM 219 331 215 49 58 10 140 RTM 211 412 191 63 30 18 142 Miranda 163 416 157 35 29 28 140 SCALE 188 191 182 32 61 17 129 JHTDB 140 225 122 33 28 23 105 SegSalt 189 645 201 51 36 13 141 Decompression CESM 661 584 689 92 101 53 513 RTM 786 622 626 124 64 108 510 Miranda 419 946 351 75 60 111 473 SCALE 610 553 567 68 140 53 450 JHTDB 376 425 243 70 59 60 330 SegSalt 592 1060 629 108 65 97 485
@@ 7.2.2. Compression ratios with the same error bounds
Compressing the datasets with the selected compressors under the same error bounds, we list all the compression ratios in Table 3 and 4. Table 3 is a comparison among the high-performance compressors (in which the compression ratio optimization targets are set for QoZ, FAZ, and HPEZ). HPEZ achieves the best compression ratio in all cases. On the SegSalt dataset, HPEZ has a $40\%$ $\sim$ $75\%$ compression ratio improvement over the second-best compressor. On the RTM, Miranda, and JHTDB datasets, HPEZ achieves 20%-45% compression ratio improvements over the second-best. On the CESM-ATM dataset, under the error bound of 1e-3, HPEZ has a compression ratio of about 2.36$\times$ as high as the second-best (SZ3.1). With these considerable improvements, we can assert that HPEZ is the best choice among high-performance compressors regarding optimizing the error-bound-fixed compression ratio.
@@
Dataset $\epsilon$ SZ 3.1 ZFP 0.5.5 QoZ 1.1 HPEZ Improve (%) RTM 1E-2 1764 62.9 2156 2701 25.3 1E-3 249 26.2 285 395 38.6 1E-4 55.3 14.3 58 71.1 22.6 Miranda 1E-2 574.6 46.6 977 1320 35.1 1E-3 168 25.6 181 258 42.5 1E-4 47.3 14.5 47.7 63.6 33.3 SegSalt 1E-2 856 59.1 1005 1484 47.7 1E-3 140.6 24.9 151 260 72.2 1E-4 38.2 14.9 35.9 61.7 61.5 SCALE 1E-2 167.3 14.5 160 186 11.2 1E-3 40.4 7.8 41.5 52.9 27.5 1E-4 14.1 4.6 13.4 15.4 9.2 JHTDB 1E-2 528.2 22.3 647 838 29.5 1E-3 73.2 9.8 77.8 101 29.8 1E-4 15.8 5 15.9 20.6 29.6 CESM-ATM 1E-2 373 18.2 263 675 81.0 1E-3 64.9 9.6 59.4 153 135.7 1E-4 22.9 5.8 21.7 38.9 69.9
@@
Dataset $\epsilon$ SPERR 0.6 FAZ TTHRESH HPEZ RTM 1E-2 2187 2695 782 2701 1E-3 440 642 71.4 395 1E-4 84.1 119 23.7 71.1 Miranda 1E-2 971.4 996.5 447 1320 1E-3 243.9 263.5 142 258 1E-4 74.5 93.6 55.1 63.6 SegSalt 1E-2 1219.4 1639.6 291 1484 1E-3 228.9 388.9 99.5 260 1E-4 61.3 117.3 28.8 61.7 SCALE 1E-2 103.5 177.9 80.0 186 1E-3 35.5 51.8 18.9 52.9 1E-4 15 16.8 8.4 15.4 JHTDB 1E-2 639.8 726 373 838 1E-3 89.3 90.7 65.1 101 1E-4 19.9 20.2 17.1 20.6 CESM-ATM 1E-2 1221 292 83.5 675 1E-3 150 77.4 20.4 153 1E-4 35 26.3 8.7 38.9
@@ 7.2.3. Compression rate-distortion
Figure 13. Rate-PSNR of HPEZ and high-ratio compressors (HPEZ’s speed is 2x-17x of high-ratio compressors).
@@ 7.2.4. Parallel data transfer test on the distributed database
Dataset Direction Inter-machine SZ3 ZFP QoZ 1.1 SPERR 0.6 FAZ TTHRESH HPEZ Improve Speed (GB/s) (%) CESM-ATM Anvil to Bebop 0.79 $\sim$0.91 1934 3221 1812 1560 1586 7752 1005 35.6 (41TB) Bebop to Anvil 0.95 $\sim$1.19 1614 2695 1553 1522 1544 8560 916 39.8 RTM Anvil to Bebop 0.58 $\sim$1.19 198 362 173 277 494 527 181 -4.8 (14TB) Bebop to Anvil 0.47 $\sim$1.04 189 524 166 296 474 560 182 -9.5 Miranda Anvil to Bebop 0.46 $\sim$1.04 49 84 44 72 87 121 39 11.3 (2TB) Bebop to Anvil 0.54 $\sim$0.82 46 117 49 71 86 120 43 6.5 SCALE-LetKF Anvil to Bebop 0.88 $\sim$0.94 873 1354 820 1037 782 2354 728 7.0 (13TB) Bebop to Anvil 1.05 $\sim$1.15 745 1181 707 1007 670 2002 624 6.8 JHTDB Anvil to Bebop 0.83 $\sim$1.15 567 826 527 645 583 835 417 20.9 (10TB) Bebop to Anvil 0.97 $\sim$1.18 486 707 473 648 574 883 366 22.7 SegSalt Anvil to Bebop 0.63 $\sim$1.18 163 289 174 221 251 393 137 15.9 (8TB) Bebop to Anvil 0.76 $\sim$1.06 167 241 153 213 265 300 132 14.0
@@ 7.2.5. Case study: decompression visualizations
As an example of the effectiveness of the HPEZ compression, in this section, we propose a case study of the compression tasks, visualizing the decompression outputs from various high-performance compressors. The example data input is the QS field (getting logarithmized in preprocessing) from the SCALE-LetKF dataset, and we compress it with HPEZ and 2 high-performance compressors: QoZ and ZFP (we omit SZ3 in this test because QoZ and SZ3 have close speeds and QoZ has better compression quality than SZ3) under similar compression ratios. The visualizations of 2-D slices from the original data and decompressed data are presented in Figure 15. In this case, among the decompression results with very close compression ratios, the decompression result of HPEZ (Figure 15 (b)) achieves the lowest data distortion with the highest PSNR (56.8). Moreover, regarding the magnified regions in Figure 15, compared to the decompression results of QoZ (Figure 15 (c), PSNR=52.7), HPEZ has better preserved the local data patterns in the original input (Figure 15 (a)). This case is an example to show the strong capability of HPEZ in providing high-quality compression results with high compression speed.
@@
Figure 15. Visualization of SCALE-QS field (logarithmized) and the decompressed data.
@@ 7.2.7. Ablation study
Figure 17 shows the rate-PSNR plots on the RTM dataset with QoZ 1.1, HPEZ, and the different accumulations of new design components between them. For example, in Figure 17 (a) representing the results of the RTM dataset, there is a curve showing the rate-distortion of QoZ 1.1, a curve showing the rate-distortion of QoZ 1.1 plus the interpolation re-ordering, a curve for the aforementioned one plus the natural cubic spline, and so on, eventually to the complete HPEZ. For the RTM, JHTDB, Miranda, and SegSalt datasets (Figure 17 (a) (c) (d) (f)), analyzing the rate-distortion curves we can easily find that the HPEZ interpolation designs, including interpolation re-ordering (Section 5.4), natural cubic spline (Section 5.2), and multi-dimensional interpolation (Section 5.3) all contribute to the improvement of rate-distortion. Additionally, block-wise interpolation tuning (Section 6.6) also plays an important role in optimizing the compression of their compression. Lastly, the effectiveness of multi-dimensional spline interpolation proved the generalization of Theorem 5.1 on diverse datasets and the integration of multiple interpolation optimization techniques.
In Figure 17 (b) and (e) corresponding to the CESM-ATM and SCALE-LetKF datasets, we can verify the effectiveness of dynamic dimension freezing (Section 6.3) and the Lorenzo predictor (Section 6.5) The dashed curve in Figure 17 (b) and (e) integrates all the interpolation designs in Section 5. Nevertheless, compared with QoZ 1.1 they have not refined the compression sufficiently. In contrast, the dynamic dimension freezing itself (the solid curves in Figure 17 (b) and the dash-dotted curve in Figure 17 (e)) has solely boosted the rate-distortion to a remarkable extent for those 2 datasets. Furthermore, comparing the solid curve and the dash-dotted curve in Figure 17 (e), leveraging the Lorenzo predictor has quite enhanced the compression quality of HPEZ in low-error-bound (i.e. high bit rates) cases.